# Emergence of Fluctuation Relations in UNO

### v1

@@ Abstract
Fluctuation theorems are generalisations of the second law that describe the relations between work, temperature, and free energy in thermodynamic processes and are used extensively in studies of irreversibility and entropy. Many experiments have verified these relations for different physical systems in the setting of thermodynamics. In this study, we observe the same behavior away from physical thermodynamics, namely for the card game UNO, by performing numerical simulations of the game. As the analog of work, we choose the number of steps one player needs to effect a transition in her deck; the other players and the remaining cards play the role of a finite, non-Markovian bath. We also compare our observation with is expected for a Markovian random walk.
@@ I Introduction
Fluctuation relations have become a mainstay in a variety of fields in science. They are associated most with stochastic thermodynamics [1, 2, 3, 4, 5, 6, 7], but also feature in fields like biophysics, optomechanics and quantum thermodynamics [8, 9, 10, 11, 12, 13]. Their significance lies in their generalization of the second law, being applicable even when systems are driven far out-of-equilibrium under non-quasistatic transformations. They describe how, despite a lack of thermodynamic constraints, strong relationships connecting average work, temperature and free energy hold. For this reason, fluctuation relations are invoked as measures of irreversibility and entropy, beyond the thermodynamic regime [12, 14, 15].
@@
Now, with this in mind, the following question emerges: if it is the case that fluctuation relations come from logical consistency, with the physics embedded in priors and assumptions – then it seems reasonable to expect some non-physics scenarios to obey certain fluctuation relations. There are plenty of possible contexts that may answer this question. For this report, we focus on games. In particular, an extensive-form scenario with moves-by-nature that is expressed as a card-shedding game [20, 21]. We are talking about the popular card game, UNO. We plot fluctuation relation graphs for various transitions in hand-sizes, in the context of this game. In particular, we found that when a specific class of cards are omitted from the game, most transitions adhere to Crooks’ fluctuation theorem.
We section the report in the following way:
  • In Section II, we review the rules for UNO. For those who are familiar with how the game works can skip this section entirely.
  • In Section III, we do two things. First, we run through the conceptual building blocks of a fluctuation relation experiment. Secondly, we show how we translate this to the game-theoretic context of UNO. We also clarify how our results are obtained in a non-trivial way.
  • In Section IV, we include the key results to be discussed later, noting standout features that should be explained.
  • In Section V, we provide explanations for the features raised in the previous section. In particular, we recontextualize thermodynamic notions (like temperature) to the game-theoretic setting.
  • Finally, in Section VI, we consolidate our key takeaways and conclude.
@@ II Reviewing the Rules of UNO
UNO (from Spanish and Italian for ‘one’), is a proprietary American card game developed in 1971 by Merle Robbins and is currently owned by Mattel, Inc. It is a shedding-type card game where players take turns to play a card onto a game pile, with the objective of losing all the cards in one’s hand. It is a derivative of a game from the 1940s known as Eights or Crazy Eights  [^1] These games are likewise a derivative of an even older card-shedding game called Mau-Mau.. These two earlier games could be played with a standard poker deck (sometimes with certain cards removed). UNO is played with a specially printed deck.
There are many variations to the game and house rules are very common. In this paper, we will go with the official rules [23].
The game consists of the deck (face-down), the gamepile (face-up), and the players’ hands (known only to the respective owner). During a player’s turn, they must choose and play a card from their hand onto the game pile, according to the card on top of the game pile. The card played must match either the same colour, or the same number/symbol as this topmost card. If this is not possible and there are no other legal moves, the player must draw one (and only one) card from the deck.
The game starts with every player being dealt a hand of seven cards, drawn randomly from the deck. The topmost card from the remaining deck is drawn onto the game pile. Once the starting player is designated, the turn order begins clockwise  [^2] There are official rules about how the dealer and the first player is chosen and so on. There are also rules about what to do when your handsize is reduced to one. These are irrelevant for this work.. The first player to lose all their cards wins.
@@
  • Normal Cards: These cards are separated into four colours (blue, green, red, and yellow). Each card has a number between 0 to 9 and there are two copies of each, except for 0 which only has one, for a total of seventy-six cards.
  • Special Cards: These cards come in the same four colours but they enforce special actions or rulings when played. These come in three types:
@@
    • The reverse card reverses the direction of the play, i.e. it inverts the turn order.
@@
  • Wildcards: These cards do not feature any specific colour or number. Rather, they allow the player to decide the next colour to be matched. The first variation of wildcard does just this, while a second variation also invokes the ruling of the +2 card, except it adds four cards instead of two. During a player’s turn, they can always play a wildcard in their hand regardless of what the game pile’s topmost card is. There are four copies of each variation, for a total of eight cards.
Of special note are the +2 cards and the +4 wildcards. We shall refer to them as plus-cards in this work. They will be shown to affect how much the game adheres to Crooks’ fluctuation relation, which we will now introduce.
@@ III.1 Key Elements of a Fluctuation Relations Experiment
We shall derive a detailed fluctuation theorem. The first such example was found by Evans and Searles [25]. Shortly later, Crooks derived his famous one, that reads [3]
@@
On the l.h.s. is the ratio between the distribution of work $W$ observed in the forward process and its associated reverse process, while the r.h.s. relates to the physics of the process. Crooks derived it for a purely deterministic (Hamiltonian) drive between two Gibbs state with inverse temperature $\beta$ and difference in free energy $\Delta F$. The same expression was then derived for the isothermal drive (Hamiltonian drive while the system exchanges energy with a thermal bath). This became the standard introductory scenario, see e.g. [5]. But detailed fluctuation theorems appear in many varieties: for instance, by conditioning on an observed value of some other parameter [26]; or independently for separate contributions to the total entropy production [27]; and many more. The version of Crooks’ theorem that we shall use is
$\frac{P_{f}(W|x\rightarrow y)}{P_{r}(-W|y\rightarrow x)}=e^{\beta(W-W_{0})}$ (2)
where $x$ is the initial (final) state of the target system for the forward (reverse) process, $y$ is the final (initial) state; $W$ designates work done on the target; $W_{0}$ is an offset for the origin of work ($\Delta F$ in thermal setting, but we won’t have this interpretation here).
In fact, any variable (typically entropy production), which can be written as the log-ratio between the statistics of the forward and the reverse process, obeys automatically the detailed fluctuation theorem [28, 17, 29]. To avoid falling into this triviality, experimental tests of detailed fluctuation theorem must define independently the variable to be measured (work) and the recipe for the reverse process. In the paradigmatic biophysical setting [8], the forward process being pulling a folded RNA strand till is straightens, the reverse process consists in pushing it; meanwhile, the work made on the system is measured as the energy invested to make the transition happen. We shall follow similar prescriptions in this paper.
@@ III.2 Getting Fluctuation Relations for UNO
Figure 2: In the forward process, Alice follows a protocol that is effective at shedding cards. Meanwhile, in the reverse process Alice does the opposite of whatever the forward protocol entails. See Section III.2.4 and Appendix A for details.
(a) UNO under official rules
(b) UNO under official rules, but with +2 cards and +4 variation wildcards removed.
Figure 3: Two sets of Crooks graphs for settings stated above. The data used to plot the graphs consist of $4\times 10^{6}$ UNO games, split evenly between the two different versions of the game and the forward and reverse processes in each of them. The lines are the Theil–Sen regression lines for each of the transitions, which is more robust to outliers than the simple linear regression [30, 31]. Data points which have a sample size of smaller than 10 have been removed. We note the oscillatory behaviour at small values of $W$.
Having reviewed the key elements of fluctuation theorem experiments, we move to translating them into our UNO setting.
III.2.1 Target System & Bath
@@
For this work’s UNO games, we can isolate one player, labelled as Alice, as the target system. The game pile, main deck and the other players qualify as a stochastic context, acting like a bath contributing randomness into Alice’s states. It is a finite bath with memory, i.e. non-Markovian, because the deck is finite and (anti-) correlated with the system (the cards comprising the bath are those that are not with Alice).
@@
In our simulations, all the other players in the bath are playing randomly, as opposed to Alice, which always play with a strategy (see Sec. III.2.4). The impact of changing the behaviour of the other players can be seen in Appendix D.
III.2.2 Initial & Final States
@@
Now, what are “Alice’s states”? For our context, we choose hand-sizes (that is, how many cards are in Alice’s hand) as what designates the states $x,y$ that parametrises the forward and reverse statistics. This designation can be seen as a coarse-graining of the exact microstates of the system (that being the exact set of cards at hand), which is standard when defining states in physical experiments [8, 32, 10].
III.2.3 Work
Work has a straightforward meaning in the context of physics. Unfortunately, in the case of UNO, this is not as obvious. For this experiment, we designate the “work” $W$ as the number of deck interactions Alice makes from $x$ to $y$ ($y$ to $x$) for the forward (reverse) process. That is, whenever she needs to draw cards, $W$ increases by 1. Whenever she places a card, $W$ increases by 1 as well.
While other choices are definitely possible, this definition parallels the fact that work is fundamentally a measure of how long the experimenter acts on the target. Normally, the work parameter fluctuates because of the stochastic environment. Likewise, in our card-shedding context, Alice’s deck interactions fluctuates based on her decisions and the randomness coming from the other players and the deck. This definition of work, however, does not allow for negative work. Thus, in order to maintain Crooks’ notation and compare $P_{F}(W)$ with $P_{R}(-W)$ instead of $P_{R}(W)$, we define work in the reverse process to be negative.
III.2.4 Forward & Reverse Process
@@
Finally, we need to define the forward and reverse processes. In a similar vein to other experiments with fluctuation relations, we characterize forward and reverse processes by the agent’s actions on the target for which the states are defined. For the context of this work, then, what defines the forward or reverse process has to do with how Alice makes decisions, changing the state of her hand. Hence, the forward process is designated as Alice playing according to a set of protocols that, statistically speaking, constitute a decent strategy at winning the game. For example, if Alice has a wildcard at hand, she will not play it unless she has no other legal moves. This helps her avoid the situation where she is forced to draw cards. The details of the protocol are given in Appendix A.
@@
If this is the forward process, what then constitutes the reverse process? We designate it as the protocol that does the converse of the forward protocol. Corresponding to the example given before, Alice, under the reverse process, plays any wildcard she has on hand as soon as she can. As with the physical experiments, the work protocol (coming from the agent or experimentalist) is reversed while being done in the same stochastic environment (in this case being the other players).
@@ III.3 Experimental Method
We perform our experiment by simulating $N=4\times 10^{6}$ UNO games on a computer and analysing the data of the results. Half of the games are played with normal UNO rules while in the other half the plus-cards (the +2 cards and +4 variation wildcards) are removed. Moreover, in half of the games Alice follows the forward algorithm, in the other half the reverse algorithm. In analysing a certain transition $x\rightarrow y$ we first obtain the probability distribution by sampling from the trajectories of Alice’s hand size in each game in the forward process games whenever this transition occurs, and vice versa for $y\rightarrow x$ in the reverse process games. We then compare the probability distributions $P_{f}(W|x\rightarrow y)$ and $P_{r}(-W|y\rightarrow x)$ to obtain the desired plots corresponding to the log of the ratio $P_{f}(W|x\rightarrow y)/P_{r}(-W|y\rightarrow x)$ against $W$. These graphs are shown in Fig. 3, one for ordinary UNO games and another for UNO games with the plus-cards removed. We see that Crooks’ relation is overall obeyed, with features to be analysed next.
@@ IV Results
Fig. 3 shows the plots of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ from the collected statistics of the game for different transitions. There are two versions of the plot, one for the official UNO game, and another one for the same game but with the plus-cards removed. Let us first highlight the some notable features observed in the figures, which we will discuss in detail in the next section.
In all cases, it could be seen that a linear relation is obeyed in the $W\gtrsim 5$ regime, with a bend at around $W\sim 5$ which gives a hockey stick character to the plots. The behavior at larger values of $W$ seems to become more erratic, but this is due to the increasingly low number of samples. Regression lines for each of the transitions are also given in the graph. It can clearly be seen that the gradients are different for each of the transitions. Notably, the gradients can be negative. We will discuss the interpretation of this in Sec. V.2.
We now note differences between Fig.3(b) and Fig.3(a), which show the effects of removing the so-called plus-cards from the game. First, the gradients for the same $x,y$ transition decrease when such cards are removed from the game. Furthermore, for the ordinary deck, there are more data points due to the extension of the support to both even and odd $W$ values and there is a small oscillatory behaviour happening in the small $W$ regime. Finally, when the plus-cards are removed, there are fewer data points in the large $W$ regime. We elaborate more on these specific points in Sec. V.3.

### v2

@@ Abstract
In the last two decades, fluctuation theorems have been proved formally and demonstrated experimentally for several variables (such as entropy production, work, or flux) and different noises causing the fluctuations (of either thermal or other origin; Markovian or non-Markovian). Here we report the observation of a detailed fluctuation relation in a statistical process outside thermodynamics and physics: the card game UNO. As the fluctuating variable, we consider the number of steps $W$ needed for one player’s deck to change from $x$ to $y$ number of cards. The other players and the remaining cards play the role of a finite non-Markovian bath. Numerical simulations of runs of the game show that $W$ obeys a fluctuation relation analogous to Crooks’ theorem. While the observed behavior shares some common features with infinite random walks, it also exhibits deviations that are clear signatures of non-Markovianity and the finiteness of the bath: Notably, the parameter corresponding to temperature depends strongly on the transition $x\rightarrow y$. Our paper contributes to extending the scope of fluctuation theorems beyond their usual thermodynamic setting.
@@ I Introduction
Fluctuation relations have become a mainstay in a variety of fields in science. Initially derived for linear response, these relations were later extended to nonlinear systems [1] and are now a key part of stochastic thermodynamics [2, 3, 4, 5, 6, 7], featuring also in fields like biophysics, optomechanics, and quantum thermodynamics [8, 9, 10, 11, 12, 13, 14]. Their significance lies in their generalization of the second law, being applicable even when systems are driven far out-of-equilibrium under nonquasistatic transformations. They describe how, despite a lack of thermodynamic constraints, strong relationships connecting average work, temperature and free energy hold. For this reason, fluctuation relations are invoked as measures of irreversibility and entropy, beyond the thermodynamic regime [13, 15, 16, 17].
@@
Now, with this in mind, the following question emerges: If it is the case that fluctuation relations come from logical consistency, with the physics embedded in priors and assumptions—then it seems reasonable to expect some non-physics scenarios to obey certain fluctuation relations. There are plenty of possible contexts that may answer this question. For this paper, we focus on games. In particular, an extensive-form scenario with moves-by-nature, expressed as a card-shedding game [22, 23]. We are talking about the popular card game, UNO. We plot fluctuation relation graphs for various transitions in hand sizes, in the context of this game. In particular, we found that when a specific class of cards are omitted from the game, most transitions adhere to Crooks’ fluctuation theorem.
We section the paper in the following way:
  1. (i)
    In Sec. II, we review the rules for UNO. For those who are familiar with how the game works can skip this section entirely.
  2. (ii)
    In Sec. III, we do two things. First, we run through the conceptual building blocks of a fluctuation relation experiment. Secondly, we show how we translate this to the game-theoretic context of UNO. We also clarify how our results are obtained in a nontrivial way.
  3. (iii)
    In Sec. IV, we include the key results to be discussed later, noting standout features that should be explained.
  4. (iv)
    In Sec. V, we provide explanations for the features raised in the previous section. In particular, we recontextualize thermodynamic notions (like temperature) to the game-theoretic setting.
  5. (v)
    Finally, in Sec. VI, we consolidate our key takeaways and conclude.
@@ II Reviewing the Rules of UNO
UNO (from Spanish and Italian for “one”), is a proprietary American card game developed in 1971 by Merle Robbins and is currently owned by Mattel, Inc. It is a shedding-type card game where players take turns to play a card onto a game pile, with the objective of losing all the cards in one’s hand. It is a derivative of a game from the 1940s known as Eights or Crazy Eights  [^1] These games are likewise a derivative of an even older card-shedding game called Mau-Mau.. These two earlier games could be played with a standard poker deck (sometimes with certain cards removed). UNO is played with a specially printed deck. There are many variations to the game and house rules are very common. In this paper, we will go with the official rules [25].
The game consists of the deck (face-down), the gamepile (face-up), and the players’ hands (known only to the respective owner). During a player’s turn, they must choose and play a card from their hand onto the game pile, according to the card on top of the game pile. The card played must match either the same color, or the same number/symbol as this topmost card. If this is not possible and there are no other legal moves, the player must draw one (and only one) card from the deck.
The game starts with every player being dealt a hand of seven cards, drawn randomly from the deck. The topmost card from the remaining deck is drawn onto the game pile. Once the starting player is designated, the turn order begins clockwise  [^2] There are official rules about how the dealer and the first player is chosen and so on. There are also rules about what to do when your handsize is reduced to one. These are irrelevant for this paper.. The first player to lose all their cards wins.
@@
  1. (i)
    Normal cards: These cards are separated into four colors (blue, green, red, and yellow). Each card has a number between 0 to 9 and there are two copies of each, except for 0, which only has one, for a total of seventy-six cards.
  2. (ii)
    Special cards: These cards come in the same four colors but they enforce special actions or rulings when played. These come in three types:
    1. (a)
@@
    2. (b)
      The reverse card reverses the direction of the play, i.e., it inverts the turn order.
    3. (c)
@@
  3. (iii)
    Wildcards: These cards do not feature any specific color or number. Rather, they allow the player to decide the next color to be matched. The first variation of wildcard does just this, while a second variation also invokes the ruling of the +2 card except that it adds four cards instead of two. During a player’s turn, they can always play a wildcard in their hand regardless of what the game pile’s topmost card is. There are four copies of each variation, for a total of eight cards.
Of special note are the +2 cards and the +4 wildcards. We shall refer to them as plus-cards in this paper. They will be shown to affect how much the game adheres to Crooks’ fluctuation relation, which we will now introduce.
@@ III.1 Key elements of a fluctuation relations experiment
The first detailed fluctuation theorem was derived for the thermodynamic work [1]. Later, similar expressions have been derived for other fluctuating variables: for instance, entropy production [6, 27] or flux in a diffusion process [28]. Our paper being about a game of cards, it is agnostic as to the physical analogy that we adopt. For definiteness, we refer to possibly the most famous detailed fluctuation theorem, derived for fluctuating work by Crooks in 1998 [3]:
@@
On the left-hand side is the ratio between the distribution of work $W$ observed in the forward process and its associated reverse process, while the right-hand side relates to the physics of the process. Crooks derived it for a purely deterministic (Hamiltonian) drive between two Gibbs state with inverse temperature $\beta$ and difference in free energy $\Delta F$. The same expression was later derived for the isothermal drive (the system undergoing evolution generated by a Hamiltonian drive while exchanging energy with a thermal bath). This became the standard introductory scenario, see e.g., [5]. Related detailed fluctuation theorems appear in many varieties: for instance, by conditioning on an observed value of some other parameter [29]. The version of Crooks’ theorem that we shall use is
$\frac{P_{f}(W|x\rightarrow y)}{P_{r}(-W|y\rightarrow x)}=e^{\beta(W-W_{0}(x,y))},$ (2)
where $x$ is the initial (final) state of the target system for the forward (reverse) process, $y$ is the final (initial) state; $W$ designates work done on the target; $W_{0}(x,y)$ is an offset for the origin of work ($\Delta F$ in thermal setting, but we would not have this interpretation here).
In fact, any variable (typically entropy production), that can be written as the log-ratio between the statistics of the forward and the reverse processes, automatically obeys the detailed fluctuation theorem [30, 19, 31]. To avoid falling into this triviality, experimental tests of detailed fluctuation theorem must define independently the variable to be measured (“work”) and the recipe for the reverse process. In the paradigmatic biophysical setting [8], the forward process is pulling a folded RNA strand until it straightens, the reverse process consists in pushing it; meanwhile, the work made on the system is measured as the energy invested to make the transition happen. We shall follow similar prescriptions in this paper.
@@ III.2 Getting fluctuation relations for UNO
Figure 2: In the forward process, Alice follows a protocol that is effective at shedding cards. Meanwhile, in the reverse process Alice does the opposite of whatever the forward protocol entails. See Sec. III.2.4 and Appendix A for details.
Figure 3: Two sets of Crooks’ graphs for (a) UNO under official rules and (b) UNO without the +2 and +4 cards. The data used to plot the graphs consist of $10\times 10^{6}$ UNO games, split evenly between the two different versions of the game and the forward and reverse processes in each of them. The lines are the Theil–Sen regression lines for each of the transitions, which is more robust to outliers than the simple linear regression [32, 33]. Variable binning is used for values with less than 100 occurrences [34]. The deviation from the straight line at small values of $W$ is relevant; that at large values is likely caused by insufficient sampling.
Having reviewed the key elements of fluctuation theorem experiments, we move to translate them into our UNO setting.
III.2.1 Target system and bath
@@
For this study’s UNO games, we can isolate one player, labeled as Alice, as the target system. The game pile, main deck, and the other players form a stochastic context that acts as a bath contributing randomness into Alice’s states. It is a finite bath with memory, i.e. non-Markovian, because the deck is finite and (anti)correlated with the system, that is, the cards comprising the bath are those that are not with Alice.
@@
In our simulations, all the other players in the bath are playing randomly, as opposed to Alice, which always play with a strategy (see Sec. III.2.4). The impact of changing the behavior of the other players can be seen in Appendix B.
III.2.2 Initial and final states
@@
What are “Alice’s states”? The most straightforward option is to use the set of the cards in Alice’s hand as the states. However, such a definition would lead to a state space that is too vast to work with as the number of possible combination of cards is extremely large. Thus, we choose the handsizes (that is, how many cards are in Alice’s hand) as the states $x,y$ that parametrises the forward and reverse statistics. This designation can be seen as a coarse graining of the exact microstates of the system (that being the exact set of cards at hand), which is standard when defining states in physical experiments [8, 35, 10].
III.2.3 Fluctuating variable (“work”)
As fluctuating variable $W$ we take the number of deck interactions Alice makes to go from $x$ to $y$ ($y$ to $x$) for the forward (reverse) process. That is, whenever she needs to draw cards, $W$ increases by 1. Whenever she places a card, $W$ increases by 1 as well. This fluctuating variable is very similar to a first passage time: we explore the connection with this well-studied notion in a Markov chain in Sec. V.1. As mentioned, we shall call this fluctuating variable “work”, motivated ultimately by a narrative connection with Crooks’ theorems. That being said, in experiments, the work performed on the system is directly related to the time taken to effect the desired transition; and this connection did inspire our choice of the definition of the fluctuating variable. Also, in a thermal setting, work fluctuates because of the stochastic environment. Likewise, in our card-shedding context, Alice’s deck interactions fluctuate based on her decisions and the randomness coming from the other players and the deck. By definition, our fluctuating variable cannot take negative values. Thus, to maintain Crooks’ notation and compare $P_{F}(W)$ with $P_{R}(-W)$ instead of $P_{R}(W)$, we simply count deck interactions in the reverse process with negative integers.
III.2.4 Forward and reverse process
Finally, we need to define the forward and reverse processes. The forward and reverse probabilities are differentiated not only by their trajectories (i.e., $x\longrightarrow y$ versus $y\longrightarrow x$), but also the reversal of the work protocol done in both processes. Note that the work protocol, sometimes also referred to as the process variable, is distinct from the work discussed in the previous section. The work protocol represents the action of the experimenter on the system and is thus deterministic and under the direct control of the experimenter. The work variable is what is measure, the stochastic fluctuating variable.
@@ III.2.4 Forward and reverse process
Therefore, in a similar vein to other experiments with fluctuation relations, we characterize forward and reverse processes by the agent’s (Alice’s) actions on the system. In our context, what defines the forward or reverse process has to do with how Alice makes decisions, changing the state of her hand. Hence, the forward process is designated as Alice play according to a set of protocols that, statistically speaking, constitute a decent strategy at winning. For example, if Alice has a wildcard on hand, she will not play it unless she has no other legal moves. This helps her avoid the situation where she is forced to draw cards.
@@
We then designate as reverse process the protocol that does the converse of the forward protocol. Under the reverse process, Alice plays any wildcard she has on hand as soon as she can. This gives a strategy more likely to lose (see Fig.2). Note that under the official rules, always drawing a card every round is a valid strategy, but for less trivial behavior we shall force Alice to always put down a card whenever possible. As with the physical experiments, the work protocol (coming from the agent or experimentalist) is reversed while being done in the same stochastic environment (in this case being the other players). Details of the protocol are given in Appendix A.
@@ III.3 Experimental Method
We perform our experiment by simulating $N=10\times 10^{6}$ UNO games on a computer and analyzing the data of the results. Half of the games are played with normal UNO rules while in the other half the plus-cards (the +2 cards and +4 variation wildcards) are removed. Moreover, in half of the games Alice follows the forward algorithm, in the other half the reverse algorithm. In analyzing a certain transition $x\rightarrow y$ we first obtain the probability distribution by sampling from the trajectories of Alice’s hand size in each game in the forward process games whenever this transition occurs, and vice versa for $y\rightarrow x$ in the reverse process games. We then compare the probability distributions $P_{f}(W|x\rightarrow y)$ and $P_{r}(-W|y\rightarrow x)$ to obtain the desired plots corresponding to the log of the ratio $P_{f}(W|x\rightarrow y)/P_{r}(-W|y\rightarrow x)$ against $W$. These graphs are shown in Fig. 3, one for ordinary UNO games and another for UNO games with the plus-cards removed. We see that Crooks’ relation is overall obeyed, with features to be analyzed next.
@@ IV Results
Figure 3 shows the plots of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ from the collected statistics of the game for different transitions. There are two versions of the plot, one for the official UNO game, and another one for the same game but with the plus-cards removed. Let us first highlight some notable features observed in the figures, which we will discuss in detail in the next section. The individual probability mass functions (pmf) of the forward and reverse processes can also be seen in Fig. 4.
In all cases, it could be seen that a linear relation is obeyed in the $W\gtrsim 5$ regime, with a bend around $W\sim 5$, which gives a hockey-stick character to the plots. The behavior at larger values of $W$ seems to become more erratic, but this is caused by the increasingly low number of samples. Regression lines for each of the transitions are also given in the graph. Let us note here that prior to this, there is no reason to expect the linear relation found here. In fact, if we had defined the work differently (for example, by the number of wild cards or red cards used in effecting a transition), we would not see this relationship. We discuss more the underlying reason for the linear relationship in Sec. V.1.
It can clearly be seen that the gradients are different for each of the transitions. Notably, the gradients can be negative. We will discuss the interpretation of this in Sec. V.2. Furthermore, there are differences between Figs. 3(b) and 3(a), which show the effects of removing the so-called plus-cards from the game. First, the gradients for the same $x,y$ transition decrease when such cards are removed from the game. Furthermore, for the ordinary deck, there are more data points due to the extension of the support to both even and odd $W$ values and there is a small oscillatory behavior happening in the small $W$ regime. Finally, when the plus-cards are removed, there are fewer data points in the large $W$ regime. We elaborate more on these specific points in Sec. V.3.